# REPS: Recycled Entropy Packet Spraying for Adaptive Load Balancing and Failure Mitigation

### v5

@@ 4.3.1. Healthy Symmetric Network Conditions
In this Section we evaluate the performance of REPS in a simple setting where there is no oversubscription and there are no failures, meaning the network is perfectly symmetrical. Intuitively, this seems the best situation for oblivious packet spraying since evenly splitting the packets across multiple links should result in the best performance. However, as we will see later in this section and based on our theoretical model in Section 5, this is not the case as REPS still offers an up to 25% advantage over OPS. This is because of ECMP collisions that still happen with OPS. While over long period of time, each link will be evenly used, there will still be short-term collisions happening that will increase and decrease the link utilization of certain links.
@@
In the same configuration with REPS, we notice a major difference in the bottom plot of Figure 1. In particular REPS converge quickly to a configuration where each queue is kept below $K_{min}$ (note that the only guarantee here is that all queues will be below $K_{min}$, not necessarily all at the same value). At the same time, we can also notice that all the ports converge to the perfect selection rate of 400 Gbps. While the overall completion time is only about 4% better than OPS, the smaller queues provide a better guarantees for system low latency traffic. Moreover, as we can see in Figure 2, this gap expands as we increase the message size.
@@
Figure 4. REPS performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives and an asymmetric network due to 2% of the TOR uplinks being offline.
@@ 4.3.2. Asymmetric Network Conditions
We observe that while OPS chooses each port equally, irrespective of its actual bandwidth, REPS eventually converges to a stable configuration where the slower uplink is used less frequently. This results in both stabler queues but, more importantly, a much faster completion time ($1400\;\mu\text{s}$ for OPS and $756\;\mu\text{s}$ for REPS).
@@
We now shift our focus to more general results when encountering asymmetries in a network. For space constraints, we focus mostly on the case where some of the links have a lower sending rate. In our first experiment we run synthetic benchmarks where $2\%$ of the TOR uplinks, chosen randomly, have been downgraded to 200 Gbps. In Figure 4, we can see results similar to before where REPS gets an up to 350% advantage over ECMP and 10% advantage over the second best algorithm (usually BitMap). In the DC traces we can see a higher difference due to the asymmetry in the network. At 100% load REPS gets a 25% advantage over the second best algorithm and a 1000% advantage over ECMP. We show the results for several AI collectives. We note how for AllToAll REPS keep a small but significant advantage and in the AllReduce a sizable 50% advantage over the second best performing algorithm.
In Figure 5, we showcase one example of REPS sharing traffic together with background ECMP traffic (we assume 10% of the traffic is ECMP). In this case, REPS: 1) shifts REPS traffic away from ECMP traffic in order to not slow down REPS traffic, 2) helps background traffic by ensuring that it will not be slowed down by REPS traffic. This also highlights the possibility of incrementally deploying REPS on ECMP-base systems. Finally, we note that WCMP (Zhou et al., 2014) could be used to enhance ECMP performance in the case of a topology with known asymmetries, but would not help as much in the case of unpredictable mixed traffic or sudden temporary asymmetries.
Figure 6. REPS vs. OPS in a 32 MiB permutation with two cables’ failure (a shorter one and a longer one).
@@ 4.3.3. Network Failures
We focus our attention to cases where the network encounters a failure during operation. We collect data from several previous works on networking failures and also study internal logs to simulate the most commonly reported cases (Gunawi et al., 2018; Singh et al., 2020, 2021). Since the probability of a failure happening during a short simulation is low, we simulate worst case scenarios where we force individual failures to happen during the runtime of our simulations. Most particularly we focus on total or partial failures of cables and switches. We note that we limit our failures to components that would not prevent the workload from completing (single point of failure).
@@
In Figure 6, OPS keeps choosing all paths equally (although at lower rate due to CC activation), while REPS, once it enters freezing mode, stops selecting the failing paths all together after only one timeout period (order of tens of microseconds). Afterwards, once the failure stops, REPS also exits freezing mode and converges once again quickly to use all paths. The overall result is that, compared to OPS, REPS completes the workload more than 35% faster even with such a short failure and, more importantly, reduces the number of dropped packets by $2.5\times$. We showcase a similar analysis but for incremental and concurrent failures in Appendix D.3.
@@
We showcase three cases with a series of failure modes in Figure 7. We introduce each failure mode after a fixed interval to highlight how REPS dynamically reacts to changing network conditions, and then let the failure persist for the remainder of the experiment. We can see when dealing with total failures that REPS provides a dramatic speedup over OPS but also other load balancer algorithms. This is because of freezing mode which helps REPS stick to the safe path configuration in its cache, allowing it to react immediately upon failure, much faster than the time ECMP routing takes to converge and exclude the failing path. Positively, we note that the gains with REPS increase with the amount of failures. Furthermore, random drops do not affect negatively REPS performance.
@@ 4.5.1. ACK Coalescing
We have primarily evaluated REPS without ACK coalescing, as this configuration allows REPS to operate with the most up to date data. However, some transport protocols permit ACK coalescing, where the receiver sends an ACK packet only after receiving $n$ data packets from the sender. In theory, the coalesced ACK packet could return all previous non-ECN marked entropies in its header, but we focus on the worst case scenario for REPS where the ACK only carries the entropy of the packet that triggered it.
Figure 11. Performance with different ACK coalescing ratios during a 8MiB permutation.
We run a real 8 MiB permutation workload simulating different ACK coalescing ratios. As we can see in the left part of Figure 11, REPS with 2:1, 4:1 and 8:1 coalescing ratios does significantly better than OPS, while with 16:1 it starts losing its advantage. However, it should be noted that in case of asymmetries or failures (right Figure 11), REPS does much better due to its various improvements over OPS even at 16:1 ratio. We confirm these results theoretically in Appendix D.1.
Finally, we note that the trade-off of sending more ACKs is worth the effort if the underlining hardware supports such rate and the impact on the network traffic is minimal (~1%) since ACKs are relatively small (64B) compared to the packet size usually used in modern interconnects (4KiB or 8KiB).
@@ 4.5.2. EVS Size
In Appendix B, we prove that OPS routing, done through EVs and ECMP hashing, requires a large EVS to work correctly and reduce collisions. Moreover, this requirement also grows with the number of output ports in a switch as demonstrated in Section 5.1. On the other hand, REPS, due to its adaptive nature, can drastically reduce the EVS size and still work well. We show this in the left plot of Figure 12 where we compare, in a real scenario, OPS and REPS when using 32, 256, and 64K EVs. REPS works equally well with 256 and 64K EVs and is only 8% slower with 32 EVs. On the other hand, OPS is 21% and 64% slower with 256 and 32 EVs when compared to 64K. This confirms that REPS could potentially work well even just with 1 byte for the EVS. We note that while OPS could be implemented without using EVS—such as through round-robin selection or by making random choices directly at the switch—these approaches introduce additional challenges (Section 5.2).
@@ 4.5.3. Different CC Algorithms
REPS has been designed to work with any CC algorithm as long as there is no over-reaction to out-of-order packets and ECN support. In this paper, we have, so far, evaluated REPS working alongside a tuned version of DCTCP. However, as we will see, REPS can work well even with other algorithms. Moreover, we envision also a version of REPS that could work just with delay if ECN is not supported but we do not go into details here. For example, in the right plot of Figure 12, we run a 8 MiB permutation workload without failure for DCTCP, EQDS and a proprietary CC algorithm. REPS can help all of these CC algorithms when compared to OPS.
Figure 12. Performance with different EVs sizes and $C$s during a 8MiB permutation.
@@ 5.1. Recycled Balls-into-Bins Model
In OPS, balls are allocated to bins uniformly at random. In what follows, we assume the EVS is sufficiently large (i.e., 16 bits), allowing us to model the assignments as uniformly random. If balls arrive at a rate of $\lambda n$ for $\lambda<1$, the process remains stable. The maximum load at any time step is $O\left(\frac{1}{1-\lambda}\log\frac{n}{1-\lambda}\right)$ with high probability (Berenbrink et al., 2018) and with probability approaching $1$, there is always a bin containing $\Omega\left(\frac{1}{1-\lambda}\log n\right)$ balls (Berenbrink et al., 2018). In the limit as $\lambda\rightarrow 1$, this implies that some bin will eventually become arbitrarily overloaded. In the context of load-balancing, this means that at the maximum injection rate, oblivious random spraying leads to unbounded queue lengths. Intuitively, this occurs because $n$ balls are introduced at each step, but fewer than $n$ may be removed, as some output ports may remain unselected.
@@
Figure 13. Simulating 10000 rounds of balls-into-bins.
@@
The proof in Appendix C shows that recycled balls-into-bins converges as bins fill and stabilize below the threshold.
@@ 6. Related Work
The literature for load balancing designs is vast, however most of it focuses on TCP-based deployments where out-of-order packets are not desirable (Benson et al., 2011; Zhang et al., 2017; Al-Fares et al., 2010; Vanini et al., 2017; Dixit et al., 2013). In contrast, REPS aligns with a new transport paradigm (UltraEthernet (Consortium, 2024), Adaptive RoCE by NVIDIA (NVIDIA, 2024), Falcon (goo, 2023)) where out-of-order packets help to leverage the capacity of multi-path networks.
Generally, most load balancers work at different granularity: per-flow, sub-flow or per-packet. ECMP (Hopps, 2009) is a per-flow solution and susceptible to flow collisions and are oblivious to congested paths. Solutions like Hedera (Al-Fares et al., 2010) and MicroTE (Benson et al., 2011) require a global controller which is not desirable in production datacenters (Gangidi et al., 2024). For these reasons, many sub-flow solutions like Flowlet Switching (Vanini et al., 2017), Presto (He et al., 2015), CONGA (Alizadeh et al., 2014), PLB (Qureshi et al., 2022), FlowBender (Kabbani et al., 2014) have been proposed. However, some of these solutions are still congestion oblivious (Flowlet, Presto), react too slowly for AI/ML bursty and intense traffic (PLB, FlowBender), require specialized switches (CONGA). Moreover, most of these solutions are unable to quickly deal with blackholes and failures.
@@
Proteus (Hu et al., 2023) focuses on optimizing the load balancing of lossless PFC (Priority Flow Control) networks while REPS focuses on lossy networks.
@@ 7. Conclusion
We presented REPS, a simple, lightweight, yet highly effective load-balancing mechanism designed to meet the constraints of next-generation datacenter networks tailored for AI workloads. As demonstrated in our extensive evaluations, conducted through both simulations and FPGA hardware, REPS’ adaptive entropy caching enhances end-to-end performance across multiple critical metrics, including average flow completion time, runtime, and packet loss. REPS outperforms ECMP and OPS by up to 6x and 1.25x in symmetric networks, and by up to 4.5x and 2x in asymmetric networks, outperforming OPS by as much as 100x during short-term transient link failures while reducing packet drops by over 70x. Moreover, REPS consistently beats or matches other state-of-the-art algorithms in all scenarios evaluated. We also showed how REPS can work well under various network configurations demonstrating its flexibility to adapt to different scenarios while remaining light-weight on NIC memory requiring only 25 bytes of state per-connection.

### v6

@@ 4.3.1. Healthy Symmetric Network Conditions
In this Section we evaluate the performance of REPS in a simple setting where there is no oversubscription and there are no failures, meaning the network is perfectly symmetrical. Intuitively, this seems the best situation for oblivious packet spraying since evenly splitting the packets across multiple links should result in the best performance. However, as we will see later in this section and based on our theoretical model in Section 5, this is not the case as REPS still offers an up to 25% advantage over OPS. This is because of ECMP collisions that still happen with OPS. While over a long period of time, each link will be evenly used, there will still be short-term collisions happening that will increase and decrease the link utilization of certain links.
@@
In the same configuration with REPS, we notice a major difference in the bottom plot of Figure 2. In particular REPS converges quickly to a configuration where each queue is kept below $K_{min}$ (note that the only guarantee here is that all queues will be below $K_{min}$, not necessarily all at the same value). At the same time, we can also notice that all the ports converge to the perfect selection rate of 400 Gbps. While the overall completion time is only about 4% better than OPS, the smaller queues provide a better guarantees for system low latency traffic. Moreover, as we can see in Figure 3, this gap expands as we increase the message size.
@@
Figure 5. REPS performance in synthetic benchmarks (I.=Incast, P.=Permutation, T.=Tornado), DC traces and AI collectives and an asymmetric network due to 3% of the TOR uplinks having downgraded bandwidth.
@@ 4.3.2. Asymmetric Network Conditions
We observe that while OPS chooses each port equally, irrespective of its actual bandwidth, REPS eventually converges to a stable configuration where the slower uplink is used less frequently. This results in both more stable queues but, more importantly, a much faster completion time ($1400\;\mu\text{s}$ for OPS and $799\;\mu\text{s}$ for REPS).
@@
We now shift our focus to more general results when encountering asymmetries in a network. For space constraints, we focus mostly on the case where some of the links have a lower sending rate. In our first experiment we run synthetic benchmarks where $3\%$ of the TOR uplinks, chosen randomly, have been downgraded to 200 Gbps. In Figure 5, we can see results similar to before where REPS gets an up to 500% advantage over ECMP and 10% advantage over the second best algorithm (usually BitMap). In the DC traces we can see a higher difference due to the asymmetry in the network. At 100% load REPS gets a 25% advantage over the second best algorithm and a 1000% advantage over ECMP. We show the results for several AI collectives. We note how for AllToAll REPS keeps a small but significant advantage and in the AllReduce a sizable 30% advantage over the second best performing algorithm.
In Figure 6, we showcase one example of REPS sharing traffic together with background ECMP traffic (we assume 10% of the traffic is ECMP). In this case, REPS: 1) shifts REPS traffic away from ECMP traffic in order to not slow down REPS traffic, 2) helps background traffic by ensuring that it will not be slowed down by REPS traffic. This also highlights the possibility of incrementally deploying REPS on ECMP-based systems. Finally, we note that WCMP (Zhou et al., 2014) could be used to enhance ECMP performance in the case of a topology with known asymmetries, but would not help as much in the case of unpredictable mixed traffic or sudden temporary asymmetries.
@@ 4.3.3. Network Failures
We focus our attention to cases where the network encounters a failure during operation. We collect data from several previous works on networking failures and also study internal logs to simulate the most commonly reported cases (Gunawi et al., 2018; Singh et al., 2020, 2021). Since the probability of a failure happening during a short simulation is low, we simulate worst case scenarios where we force individual failures to happen during the runtime of our simulations. In particular, we focus on total or partial failures of cables and switches. We note that we limit our failures to components that would not prevent the workload from completing (single point of failure).
Figure 7. REPS vs. OPS in a 64 MiB permutation with two cable failures (a shorter one and a longer one).
@@
In Figure 7, OPS keeps choosing all paths equally (although at lower rate due to CC activation), while REPS, once it enters freezing mode, stops selecting the failing paths all together after only one timeout period (order of tens of microseconds). Afterwards, once the failure stops, REPS also exits freezing mode and converges once again to use all paths. The overall result is that, compared to OPS, REPS completes the workload more than 35% faster even with such a short failure and, more importantly, reduces the number of dropped packets by $2.5\times$. We showcase a similar analysis for incremental and concurrent failures in Appendix C.3.
@@
We showcase three cases with a series of failure modes in Figure 8. We introduce each failure mode after a fixed interval to highlight how REPS dynamically reacts to changing network conditions, and then let the failure persist for the remainder of the experiment. We can see when dealing with total failures that REPS provides a dramatic speedup over OPS but also other load balancer algorithms. This is because of freezing mode which helps REPS stick to the safe path configuration in its cache, allowing it to react immediately upon failure, much faster than the time ECMP routing takes to converge and exclude the failing path. Positively, we note that the gains with REPS increase with the amount of failures. Furthermore, random drops do not affect negatively REPS performance. MPRDMA also performs well on average thanks to its self-clocking mechanism.
@@ 4.5.1. ACK Coalescing
We have primarily evaluated REPS without ACK coalescing, as this configuration allows REPS to operate with the most up to date data. However, some transport protocols permit ACK coalescing, where the receiver sends an ACK packet only after receiving $n$ data packets from the sender. In theory, the coalesced ACK packet could return all previous non-ECN marked entropies in its header (we call this configuration ACK+Carry EVs). As an alternative approach, each entropy could be assigned a lifespan in the REPS buffer and reused $n$ times (ACK+Reuse EVs).
Figure 12. Performance with different ACK coalescing ratios during an 8MiB permutation.
We run an 8 MiB permutation simulating different ACK coalescing ratios and only the standard REPS configuration. As we can see in the left part of Figure 12, REPS with 2:1, 4:1, and 8:1 coalescing ratios does significantly better than OPS, while with 16:1 it starts losing its advantage. However, it should be noted that in case of asymmetries or failures (right Figure 12), REPS does much better even at 16:1 ratio. We confirm these results theoretically in Appendix C.1.
Figure 13 demonstrates that Carry EVs and Reuse EVs are the preferred REPS variants for ACK coalescing, particularly under high coalescing ratios. We leave extensions of this approach to future work.
Figure 13. Different REPS variants for ACK coalescing.
Finally, we note that the trade-off of sending more ACKs is worth the effort if the underlying hardware supports such rate and the impact on the network traffic is minimal (~1%).
@@ 4.5.2. EVS Size
Practical implementations make use of hashing functions to map EVs onto output ports. We show, first theoretically and then with simulations, that while using a small EVS can lead to high inherent load balancing issues, using $2^{16}$ EVs is very close to uniformly random.
We use a balls-into-bins model (Raab and Steger, 1998), where $m$ balls are thrown, uniformly and independently, at random into $n$ bins. The goal is to determine the largest number of balls in any bin, referred to as the maximum load $l(m,n)$. We define the load imbalance $\lambda_{m,n}$ as $\frac{l(m,n)}{m/n}-1$, representing the extent to which the most heavily loaded bin exceeds the average.
We confirm this theoretical analysis with simulations of the load imbalance. Figures 14(a) and 14(b) show the distribution of the load imbalance for $1$ and $32$ unique flows, respectively. We note that each flow is from a different sender and will hence have different header fields that will be used in the hashing function regardless of the EV value. We note that for each case, we throw for each active flow a number of balls equal to the EVS size, with each ball being a unique EV. We can see that for 32 flows, choosing less than $2^{8}$ EVs can lead to more than $10\%$ load-imbalance, whereas $2^{16}$ EVs guarantee less than a $1\%$ load imbalance.
((a)) 1 flow active
((b)) 32 flows active
Figure 14. Expected load imbalance at a switch (32 uplinks).
Moreover, this requirement also grows with the number of output ports in a switch as demonstrated in Section 5.1. On the other hand, REPS, due to its adaptive nature, can drastically reduce the EVS size and still work well. We show this in the left plot of Figure 15 where we compare, in a real scenario, OPS and REPS when using 32, 256, and 64K EVs. REPS works equally well with 256 and 64K EVs and is only 8% slower with 32 EVs. On the other hand, OPS is 21% and 64% slower with 256 and 32 EVs when compared to 64K. This confirms that REPS could potentially work well even just with 1 byte for the EVS. We note that while OPS could be implemented without using EVS, such as through round-robin selection or by making random choices directly at the switch, these approaches introduce additional challenges (Section 5.2).
@@ 4.5.3. Different CC Algorithms
REPS has been designed to work with any CC algorithm as long as there is no over-reaction to out-of-order packets and ECN support. In this paper, we have, so far, evaluated REPS working alongside a tuned version of DCTCP (similar to what is used in MPRDMA (Lu et al., 2018)). However, as we will see, REPS can work well even with other algorithms. Moreover, we envision also a version of REPS that could work just with delay if ECN is not supported but we do not go into details here. For example, in the right plot of Figure 15, we run an 8 MiB permutation workload without failure for DCTCP, EQDS and a proprietary CC algorithm. REPS can help all of these CC algorithms when compared to OPS.
Intuitively, REPS works on a per-packet basis, meaning each flow simultaneously uses many paths. If one path starts encountering congestion, REPS will quickly route away from it. Temporarily, the CC algorithm might slightly lower the congestion window or sending rate. For instance, if it uses only ECN as its signal and only one path is congested, the decrease would be minimal (assuming a CC that does not overreact to each ECN-mark). Moreover, non-ECN packets from uncongested paths would quickly restore the window. If all paths are congested, REPS cannot add capacity, and the CC rightly lowers the congestion window; this is precisely what we observe in the incast experiments where every balancer behaves the same (Figure 3). In these cases REPS (like all routing algorithms) cannot fix the issue and the CC must kick-in to correctly adjust the sending rate. Finally, we note that one could potentially optimize the CC further for a smoother REPS integration. For example, the CC could activate only after a certain degree of congestion is detected, while leaving the initial work to the load balancer. Having said that, this is mostly orthogonal to REPS and also affects other load balancers.
Figure 15. Performance with different EVs sizes and $C$s during an 8MiB permutation.
4.5.4. Topology Scaling
We then shift to verifying REPS performance when changing the scale of the topology. To do this, we run a 8MiB tornado workload and vary the number of nodes in the topology. We range from 128 nodes (switches with 16 ports) to 8192 nodes (switches with 128 ports). Moreover, we also vary the EVS size and compare once again REPS to OPS. Figure 16 shows the maximum completion time for each configuration and we run each experiment multiple times to account for randomness in the initial EVs. As we can see, REPS performs well for all topology sizes and almost all EVS sizes with only a small regression when using 16 EVs. This, potentially, means that REPS could work with far less than 16 bits needed for the EV, even at large scale. On the other hand, OPS suffers much more due to its non-adaptivity when having a limited number of EVs with 16 values running more than twice slower than when using the full 16 bits (65536 possible values). We also observe a slight upward trend in the maximum FCT with OPS as topology size increases. The key point is that, with REPS, even if a limited number of EVs prevents each flow from using all available paths, the increased adaptivity compensates for this. In addition, other packet-header fields contribute additional randomness because switch hashing functions combine them with the EV. This observation is consistent with the analysis in Section 4.5.2.
Figure 16. Performance with different EVS sizes and topology sizes during an 8MiB permutation.
@@ 5.1. Recycled Balls-into-Bins Model
In OPS, balls are allocated to bins uniformly at random. In what follows, we assume the EVS is sufficiently large (i.e., 16 bits), allowing us to model the assignments as uniformly random. If balls arrive at a rate of $\lambda n$ for $\lambda<1$, the process remains stable. The maximum load at any time step is $O\left(\frac{1}{1-\lambda}\log\frac{n}{1-\lambda}\right)$ with high probability (Berenbrink et al., 2018) and with probability approaching $1$, there is always a bin containing $\Omega\left(\frac{1}{1-\lambda}\log n\right)$ balls (Berenbrink et al., 2018). In the limit as $\lambda\rightarrow 1$, this implies that some bin will eventually become arbitrarily overloaded. In the context of load balancing, this means that at the maximum injection rate, oblivious random spraying leads to unbounded queue lengths. Intuitively, this occurs because $n$ balls are introduced at each step, but fewer than $n$ may be removed, as some output ports may remain unselected.
@@
Figure 17. Simulating 1000 rounds of balls-into-bins.
@@
The proof in Appendix B shows that recycled balls-into-bins converges as bins fill and stabilize below the threshold.
@@ 6. Related Work
The literature for load balancing designs is vast, however most of it focuses on TCP-based deployments where out-of-order packets are not desirable (Benson et al., 2011; Zhang et al., 2017; Al-Fares et al., 2010; Vanini et al., 2017; Dixit et al., 2013). In contrast, REPS aligns with a new transport paradigm (UltraEthernet (Consortium, 2024), Adaptive RoCE by NVIDIA (NVIDIA, 2024), Falcon (Singhvi et al., 2025)) where out-of-order packets help to leverage the capacity of multi-path networks.
Generally, most load balancers work at different granularity: per-flow, sub-flow or per-packet. ECMP (Hopps, 2009) is a per-flow solution and susceptible to flow collisions and are oblivious to congested paths. Solutions like Hedera (Al-Fares et al., 2010) and MicroTE (Benson et al., 2011) require a global controller which is not desirable in production datacenters (Gangidi et al., 2024). For these reasons, many sub-flow solutions like Flowlet Switching (Vanini et al., 2017), Flowcut Switching (Bonato et al., 2026), Presto (He et al., 2015), CONGA (Alizadeh et al., 2014), PLB (Qureshi et al., 2022), FlowBender (Kabbani et al., 2014) have been proposed. However, some of these solutions are still congestion oblivious (Flowlet, Presto), react too slowly for AI/ML bursty and intense traffic (PLB, FlowBender), require specialized switches (CONGA). Moreover, most of these solutions are unable to quickly deal with blackholes and failures.
@@
Proteus (Hu et al., 2023) focuses on optimizing the load balancing of lossless PFC networks while REPS focuses on lossy networks.
@@ 7. Conclusion
We presented REPS, a simple, lightweight, yet highly effective load-balancing mechanism designed to meet the constraints of next-generation datacenter networks tailored for AI workloads. As demonstrated in our extensive evaluations, conducted through both simulations and FPGA hardware, REPS’ adaptive entropy caching enhances end-to-end performance across multiple critical metrics, including average flow completion time, runtime, and packet loss. REPS outperforms ECMP and OPS by up to 6x and 1.25x in symmetric networks, and by up to 5x and 2x in asymmetric networks, outperforming OPS by as much as 100x during short-term transient link failures while reducing packet drops by over 70x. Moreover, REPS consistently beats or matches other state-of-the-art algorithms in all scenarios evaluated. We also showed how REPS can work well under various network configurations demonstrating its flexibility to adapt to different scenarios while remaining light-weight on NIC memory requiring only 25 bytes of state per-connection.